OpenStudy (anonymous):

what's the limit of lim (x-5)/(x^2 -25) x-> 5+

OpenStudy (anonymous):

x^2-25=(x+5)(x-5) so you can cancel x-5 this will remain: lim 1/(x+5)=1/(5+5)=1/10

OpenStudy (anonymous):

is positive infinity wrong?

OpenStudy (anonymous):

It is, the limit is 1/10

OpenStudy (anonymous):

but when you graph it, the line approaching x=5 never ends....

OpenStudy (anonymous):

well I did not graph anything just used algebra